# Evolution of Shock Structures and QPOs After Halting BHL Accretion onto Kerr Black Hole

### v1

@@ 5 Effect of Spin and Asymptotic Velocity on LFQPOs and HFQPOs
HFQPOs, on the other hand, originate from the orbital motion of matter. More specifically, the radial epicyclic motion of matter contributes to these frequencies. Since the motion of matter is influenced by the black hole’s spin, a faster-spinning black hole results in higher HFQPO values. As confirmed numerically in Table 2, as the black hole spin increases from $a/M=0.0$ to $a/M=0.9$, the QPO frequency increases from $36$ Hz to $40.84$ Hz. Furthermore, as seen in Table 2, HFQPOs can also generate new frequencies through nonlinear couplings. The presence of nonlinear couplings indicates complex interactions between different modes of disk oscillations. Table 2 shows that for high-spin cases ( $a/M=0.9$), nonlinear couplings appear, but they vanish for moderate or zero spin. This suggests that nonlinear interactions in the inner accretion disk are strongly influenced by the black hole spin and the accretion geometry and it is also consistent with the existence literature 2013PhRvL.111d1101S ; 2024PhRvD.109j1503R .
@@
Table 2: The QPO frequencies obtained from numerical simulations, their occurrence based on the black hole’s spin parameter and asymptotic velocity, and the possible nonlinear coupling probabilities are presented.
@@
$0.5$ $0.2$ $4$, $16.8$ $38$ $No$
$0.0$ $0.2$ $No$ $36$ $No$
@@ 6 Comparison of Numerical QPO Frequencies with Observational Data
We compare the QPO frequencies obtained from the analysis of the precession motion of plasma in the strong gravitational field after the BHL accretion is halted, as given in Table 2, with the observational results presented in Table 3. In our numerical calculations, we assumed the black hole mass to be $M=10M_{\odot}$ when computing QPO frequencies in Hertz. To scale the numerically obtained frequencies according to the black hole masses provided in Table 3, we can use the formula $\nu_{new}=\nu_{old}\;\frac{10M_{\odot}}{M_{BH}}$. Using this formula, we can also recalculate our numerical frequencies for massive black holes and make a direct comparison with any given sources. For the rapidly spinning black hole case ($a/M=0.9$), our low-frequency numerical results show agreement with the observed $Type-C$ QPOs of $GRS1915+105$ and $XTEJ1550-564$. However, the numerically obtained frequencies of $36$ Hz, $38$ Hz, $40.8$ Hz, and $44.6$ Hz show consistency with the observed HFQPO frequencies of $GRS1915+105$. Furthermore, we can compare our findings with the AGN source $REJ1034+396$, which is known to host a massive black hole. In this case, if we rescale our numerical results for a supermassive black hole (SMBH) using: $\nu_{SMBH}=\frac{36}{10^{5}}=3.6x10^{-4}Hz$ and $\nu_{SMBH}=\frac{40.8}{10^{5}}=4.08x10^{-4}Hz$. The rescaled frequencies align well with the observed frequency given in Table 3 for $REJ1034+396$.
Overall, the numerical results indicate that LFQPOs only emerge in the case of supersonic accretion when the black hole has a high spin parameter. This finding is in agreement with the numerical simulations and observations Ingram2009MNRAS ; 2015MNRAS.452.3451C ; 2020MNRAS.492..804M ; 2022MNRAS.510..807S . In other words, strong LFQPOs are found to form in the plasma surrounding rapidly rotating black holes, which is also consistent with observational data. For HFQPOs, the resonance structure ($3:2$, $2:1$) observed in the numerical simulations due to nonlinear couplings perfectly matches the observational results 2015A&A…581A..35B ; 2022JCAP…08..034B . However, some of the HFQPO frequencies observed in certain sources (e.g. $300$ Hz, $450$ Hz in $GROJ1655-40$) do not appear in the numerical simulations. This suggests that additional physical effects may be necessary for their formation. Possible mechanisms that could lead to the emergence of these higher frequencies include disk resonance effects, magnetic field influences, modifications to general relativity (alternative gravity theories) , or additional strong-field GR effects. If these factors play a role in QPO formation, they might explain the presence of the higher-frequency HFQPOs seen in observations.
@@
$GRS1915+105$ 1218121812–1812 – 18 1101101–101 – 10 $41,67$ Strohmayer2001ApJ ; Misra2004 ; Liu_2021 ; Majumder_2022 ; Motta2023
$GROJ1655-40$ $6.3$ 0.1100.1100.1–100.1 – 10 $300,450$ Remillard1999ApJ ; 2012MNRAS.427..595M ; Motta2014MNRAS
$XTEJ1550-564$ $9.1$ 0.1100.1100.1–100.1 – 10 $184,276$ Varniere_2018
$H1743-322$ 8108108–108 – 10 0.1100.1100.1–100.1 – 10 $240$ 2006ApJ…637.1002R ; 2012ApJ…754L..23A
@@ 7 New Insights into QPO Formation: A Comparison with Earlier Findings
The results we obtained here are important for comparing the QPO frequencies generated by the oscillation of the shock cone around the Kerr black hole and the modes trapped inside the cone with BHL accretion. Previously, in Donmez6 , the dynamic structure of the shock cone formed around the Kerr black hole was examined under both sonic and supersonic flow conditions, depending on different black hole spin parameters. The QPO frequencies obtained are provided in Table 2 of Donmez6 . As seen in this table, only HFQPOs are formed. In contrast, in this study, by halting BHL accretion, a new plasma structure around the black hole has formed, resulting in both LFQPOs and HFQPOs. When the results of these two studies are compared with the sources given in Section 6, the QPO frequencies obtained in this study are consistent because they explain both the LFQPO and HFQPO frequencies. Therefore, the plasma structure that forms in the strong gravitational field around the black hole because of the cessation of the BHL mechanism sheds light on the physical mechanisms that can be proposed for QPOs in the literature. On the other hand, the fundamental HFQPO modes and their nonlinear coupling-generated frequencies in Table 2 of Donmez6 , as well as the increase in these frequencies with the black hole’s spin, are in perfect agreement with the results obtained in this study. In conclusion, for QPO formation in observed sources, the best physical model that can be proposed is the plasma structure found in this study. The behavior of this plasma, influenced by both precession motion and strong rotation, provides a direct match with observations, ensuring the formation of QPOs.
The decrease in the rest-mass density of matter (plasma) around the black hole and the change in its dynamic structure can affect the trapping or excitation of waves that may occur within the plasma. At the same time, it can alter the nonlinear coupling mechanism, leading to the formation of weaker and fewer QPOs in this scenario 2008MNRAS.386.2297F . However, because of these plasma changes, the variation in the density of the Comptonizing cloud alters the cooling and infall timescales, which in turn can contribute to the formation of LFQPOs 2020MNRAS.492..804M . When comparing the numerical PSD analyzes obtained in Section 4 with those of Donmez6 ; Donmez4 , it is observed that the resulting QPO frequencies are fewer in number and have lower amplitudes. Furthermore, these plasma variations have caused changes in the resonant states, frequencies, and their amplitudes 2004PASJ…56..931K . As a result, the numerical findings obtained are consistent with those in the literature and provide an explanation for the HFQPO and LFQPO observed from different sources.
@@ 8 Conclusions
In this paper, we model the complete disappearance of the shock cone formed around the black hole. The shock cone is a physical mechanism previously established in strong gravitational fields around the black holes using the BHL mechanism in both general relativity and alternative gravity theories. We examine how this shock cone mechanism evolves after the cessation of BHL accretion and observe the formation process of the newly emerging plasma structure. Eventually, after a long dynamical evolution, the matter in plasma reaches a steady-state configuration around the black hole, forming a quasi-periodically oscillating plasma structure. We investigate how the dynamic evolution of the plasma structure and the resulting precession motion depend on the black hole’s spin parameter and asymptotic velocity. Finally, by performing a power spectrum analysis, we identify the QPO frequencies that emerge from the system.
@@
It has been observed that the asymptotic velocity used in modeling the shock cone around a black hole via BHL accretion affects the plasma structure formed after BHL is halted, and consequently, the resulting QPOs. For a rapidly rotating black hole with $a/M=0.9$ and an asymptotic velocity of $V_{\infty}/c=0.1$, meaning that the flow is sonic, LFQPOs are not observed. However, LFQPOs appeared at other asymptotic velocities. Therefore, for LFQPOs to be observable, the Lense-Thirring effect, which arises solely from the black hole’s rapid rotation and its warping of spacetime, would not be sufficient. Additionally, it is known that indirectly the velocity of the gas and, consequently, its angular momentum contribute to the formation and regulation of QPOs Mukhopadhyay_2009 ; 2009MNRAS.398.1105R . However, based on the numerical simulations, it has been found for the first time that if the plasma exhibits supersonic behavior, LFQPOs emerge.
@@
As seen in numerical simulations and Table 2, in addition to the fundamental HFQPOs, other HFQPOs have been observed that generate different resonance modes (e.g. $3:2$, and $2:1$), which are consistent with observations. These frequencies result from the fundamental mode itself and the nonlinear couplings that form between subsequently generated QPOs. Furthermore, as shown in Table 2, these couplings occur only when the black hole is rapidly rotating. Thus, the presence of nonlinear couplings in rapidly rotating black holes suggests that additional oscillatory modes are introduced due to the strong frame-dragging effect and the differential rotation of the accretion disk.
The results obtained here show consistency with the observations. In microquasars $GRS1915+105$, $XTEJ1550-564$, and $GROJ1655-40$, the observed QPO frequencies range between $30-200$ Hz. The numerical simulations presented in this study yield HFQPO frequencies in the range of $36-49$ Hz, which are consistent with black holes of approximately stellar mass $M\sim 10M_{\odot}$. Additionally, the fact that the numerical simulations show double and triple the fundamental frequency suggests that the $3:2$ and $2:1$ resonance states observed in black hole QPOs can be explained through these results. This supports the idea that QPOs in black hole systems may arise from parametric resonances in the accretion disk. On the other hand, the observed frequency range for LFQPOs in black hole $X-$ray binaries is 0.1300.1300.1–300.1 – 30 Hz, whereas the numerical simulations in this study produced LFQPO frequencies in the range of $1.5-22.6$ Hz, which aligns well with observations. However, LFQPOs did not appear when $V_{\infty}/c=0.1$ for $a/M=0.9$. This indicates that LFQPOs depend on the accretion state, reinforcing their consistency with observational data.
@@
We will continue working in the future to uncover the effects of the black hole’s spin parameter, accretion dynamics, and the physical parameters governing these processes, as well as the behavior of plasma in a strong gravitational field. In upcoming studies, we will utilize modified gravity theories to investigate the influence of these alternative gravitational models on the QPO frequencies that we have identified in this study. Through this approach, we aim to determine the impact of scalar field, which represents the existence of dark matter, on the results obtained here.

### v2

@@ 5 Effect of Spin and Asymptotic Velocity on LFQPOs and HFQPOs
HFQPOs, on the other hand, originate from the orbital motion of matter. More specifically, the radial epicyclic motion of matter contributes to these frequencies. Since the motion of matter is influenced by the black hole spin, a faster-spinning black hole results in higher HFQPO values. As confirmed numerically in Table 2, as the black hole spin increases from $a/M=0.0$ to $a/M=0.9$, the QPO frequency increases from $36$ Hz to $40.84$ Hz. Furthermore, as seen in Table 2, HFQPOs can also generate new frequencies through nonlinear couplings. The presence of nonlinear couplings indicates complex interactions between different modes of disk oscillations. Table 2 shows that for high-spin cases ( $a/M=0.9$), nonlinear couplings appear, but they vanish for moderate or zero spin. This suggests that nonlinear interactions in the inner accretion disk are strongly influenced by the black hole spin and the accretion geometry and it is also consistent with the existence literature 2013PhRvL.111d1101S ; 2024PhRvD.109j1503R .
@@
Table 2: The QPO frequencies obtained from numerical simulations, their occurrence based on the black hole spin parameter and asymptotic velocity, and the possible nonlinear coupling probabilities are presented.
@@
$0.5$ $0.2$ 44, $16.8$ $38$ NoNo
$0.0$ $0.2$ NoNo $36$ NoNo
As shown in Table 2, the QPO frequencies correspond to global modes that are independent of the chosen time window of the data, can be identified throughout the disk, and are not sensitive to the specific diagnostic parameter used. This confirms that the QPOs observed in our simulations are not numerical artifacts but are instead the natural outcome of excited oscillation modes in the disk. The first test we performed is to divide the accretion rate signal, after the BHL accretion is halted and the disk reached a quasi-steady state, into multiple windows, and recompute the PSD in each segment. In all cases, almost the same frequencies reappeared. A more detailed version of this test was presented in our earlier work (2024EPJC…84..524D, ; Donmez2023arXiv231013847D, ; Donmez2024Submitted, ). In all of our numerical analyses, the power spectrum peaks are found at nearly identical frequencies. The second test, frequently applied in our analysis, is to compute the PSD at different radial locations in the flow (e.g., $r=2.3M$, $r=6.1M$, and $r=12M$). As seen in previous references (Donmez6, ; Donmez2024Univ, ; Donmez2024arXiv240701478D, ; Donmez2024arXiv240216707D, ) and confirmed again in this study, the peaks appear at nearly the same frequencies regardless of radius, with only their amplitudes varying depending on the plasma density near the black hole. This indicates that the peaks originate from global disk instabilities rather than local turbulence, demonstrating that the QPOs are global rather than local in nature (Donmez6, ; 2011MNRAS.417.2899Z, ; Donmez5, ). In addition, beyond calculating the PSD from the mass accretion rate as we did throughout this paper, we also performed the analysis using the mass density at a fixed location near the black hole, sampled at different times. We found that the resulting frequencies are essentially identical to those obtained from the accretion rate. This further confirms that the frequencies are not artifacts of the numerical method. Together, these tests demonstrate that the frequencies reported in Table 2 are robust against variations in temporal sampling, radial location, and diagnostic quantity. This strengthens the reliability of our numerical results and places them on firmer ground for comparison with observed QPOs in X-ray binaries (Remillard2006ARA&A, ; Ingram2009MNRAS, ).
@@ 6 Comparison of Numerical QPO Frequencies with Observational Data
We compare the QPO frequencies obtained from the analysis of the precession motion of plasma in the strong gravitational field after the BHL accretion is halted, as given in Table 2, with the observational results presented in Table 3. As noted in the final paragraph of the Introduction, in this paper we adopt geometrized units. In geometrized units, the frequency has units of $1/M$, where $M$ is the black hole mass. Since the frequencies we compute from the PSD analyses depend explicitly on the black hole mass, the results obtained here can be used not only to explain the QPO frequencies observed around stellar-mass black holes but also to shed light on the observational findings for supermassive black hole sources. Therefore, in order to compare the numerical QPO frequencies given in Table 2 with the observational results summarized in Table 3, we must convert our numerical results from geometrized units ($1/M$) to SI units (Hz). The conversion from geometrized to SI units is performed using $1/M=c^{3}/(GM)$ where $c$ is the speed of light and $G$ is the gravitational constant. As shown in Table 3, since our comparisons in this study are mainly with QPO frequencies observed around stellar-mass black holes and the typical masses of these sources are around $M=10M_{\odot}$, we adopt $M=10M_{\odot}$ in carrying out this conversion. In this case, the conversion formula from geometrized to SI units becomes,
$\nu_{new}=\nu_{old}\;\frac{10M_{\odot}}{M_{BH}}.$ (8)
@@
Using Eq.8, we can also recalculate our numerical frequencies for massive black holes and make a direct comparison with any given sources. For the rapidly spinning black hole case ($a/M=0.9$), our low-frequency numerical results show agreement with the observed TypeCType-C QPOs of $GRS1915+105$ and XTEJ1550564XTEJ1550-564. However, the numerically obtained frequencies of $36$ Hz, $38$ Hz, $40.8$ Hz, and $44.6$ Hz show consistency with the observed HFQPO frequencies of $GRS1915+105$. Furthermore, we can compare our findings with the AGN source $REJ1034+396$, which is known to host a massive black hole. In this case, if we rescale the frequencies found in the PSD analyses for the high-mass source $REJ1034+396$ (with a mass of $10^{6}M_{\odot}$) given in Table 3, we obtain $\nu_{SMBH}=\frac{36}{10^{5}}=3.6x10^{-4}$ Hz and $\nu_{SMBH}=\frac{40.8}{10^{5}}=4.08x10^{-4}Hz$. The rescaled frequencies align well with the observed frequency given in Table 3 for $REJ1034+396$.
Overall, the numerical results indicate that LFQPOs only emerge in the case of supersonic accretion when the black hole has a high spin parameter. This finding is in agreement with the numerical simulations and observations Ingram2009MNRAS ; 2015MNRAS.452.3451C ; 2020MNRAS.492..804M ; 2022MNRAS.510..807S . In other words, strong LFQPOs are found to form in the plasma surrounding rapidly rotating black holes, which is also consistent with observational data. For HFQPOs, the resonance structure ($3:2$, $2:1$) observed in the numerical simulations due to nonlinear couplings perfectly matches the observational results 2015A&A…581A..35B ; 2022JCAP…08..034B . However, some of the HFQPO frequencies observed in certain sources (e.g. $300$ Hz, $450$ Hz in GROJ165540GROJ1655-40) do not appear in the numerical simulations. One reason for this may be the assumptions and simplifications made in setting up the simulations. In this study, we neglect the effects of magnetic fields and numerically solve the $2D$ GRHD equations to model the plasma structure formed by matter acting as an ideal gas around the black hole, along with the resulting shock waves. Under these assumptions, LFQPOs and some HFQPOs can indeed be recovered numerically, but certain frequencies reported in observations do not appear in the simulations. In contrast, the higher-frequency QPOs (e.g., 300450300–450 Hz) may require additional physical mechanisms not captured here, such as magneto-rotational turbulence, disk inhomogeneities, or strong magnetic stresses. These processes can enhance the coupling between the inner disk and the black hole spin. Furthermore, our analysis assumes a black hole mass of $M=10M_{\odot}$, and while scaling relations allow comparisons across systems, uncertainties in the true mass and spin of GROJ165540GROJ1655-40 introduce further shifts in the predicted frequencies. Furthermore, because the accretion mechanism in this study is modeled in the equatorial plane, the vertical epicyclic frequencies are neglected, which may also explain why the $300$ and $450$ Hz features do not emerge in our simulations. Therefore, the absence of the $300$ Hz and $450$ Hz frequencies in our results does not imply that the mechanism modeled here is incapable of producing them, but rather that the restricted hydrodynamic setup isolates only a subset of the physically relevant oscillatory modes. A complete $3D$ GRMHD treatment of the problem considered in this paper would allow us to investigate whether the mechanisms mentioned above contribute to the formation of these HFQPOs. Lastly, in recent years, alternative theories of gravity have been proposed that may produce results capable of explaining the origin of these frequencies. If such a scenario holds, these HFQPOs could serve as important observational probes for testing strong gravity.
@@
$GRS1915+105$ 121812–18 1101–10 $41,67$ Strohmayer2001ApJ ; Misra2004 ; Liu_2021 ; Majumder_2022 ; Motta2023
GROJ165540GROJ1655-40 $6.3$ 0.1100.1–10 $300,450$ Remillard1999ApJ ; 2012MNRAS.427..595M ; Motta2014MNRAS
XTEJ1550564XTEJ1550-564 $9.1$ 0.1100.1–10 $184,276$ Varniere_2018
H1743322H1743-322 8108–10 0.1100.1–10 $240$ 2006ApJ…637.1002R ; 2012ApJ…754L..23A
@@ 7 New Insights into QPO Formation: A Comparison with Earlier Findings
The results we obtained here are important for comparing the QPO frequencies generated by the oscillation of the shock cone around the Kerr black hole and the modes trapped inside the cone with BHL accretion. Previously, in Donmez6 , the dynamic structure of the shock cone formed around the Kerr black hole was examined under both sonic and supersonic flow conditions, depending on different black hole spin parameters. The QPO frequencies obtained are provided in Table 2 of Donmez6 . As seen in this table, only HFQPOs are formed. In contrast, in this study, by halting BHL accretion, a new plasma structure around the black hole has formed, resulting in both LFQPOs and HFQPOs. When the results of these two studies are compared with the sources given in Section 6, the QPO frequencies obtained in this study are consistent because they explain both the LFQPO and HFQPO frequencies. Therefore, the plasma structure that forms in the strong gravitational field around the black hole because of the cessation of the BHL mechanism sheds light on the physical mechanisms that can be proposed for QPOs in the literature. On the other hand, the fundamental HFQPO modes and their nonlinear coupling-generated frequencies in Table 2 of Donmez6 , as well as the increase in these frequencies with the black hole spin, are in perfect agreement with the results obtained in this study. In conclusion, for QPO formation in observed sources, the best physical model that can be proposed is the plasma structure found in this study. The behavior of this plasma, influenced by both precession motion and strong rotation, provides a direct match with observations, ensuring the formation of QPOs.
The decrease in the rest-mass density of matter (plasma) around the black hole and the change in its dynamic structure can affect the trapping or excitation of waves that may occur within the plasma. At the same time, it can alter the nonlinear coupling mechanism, leading to the formation of weaker and fewer QPOs in this scenario 2008MNRAS.386.2297F . However, because of these plasma changes, the variation in the density of the Comptonizing cloud alters the cooling and infall timescales, which in turn can contribute to the formation of LFQPOs 2020MNRAS.492..804M . When comparing the numerical PSD analyzes obtained in Section 4 with those of Donmez6 ; Donmez4 , it is observed that the resulting QPO frequencies are smaller in number and have lower amplitudes. Furthermore, these plasma variations have caused changes in the resonant states, frequencies, and their amplitudes 2004PASJ…56..931K . As a result, the numerical findings obtained are consistent with those in the literature and provide an explanation for the HFQPO and LFQPO observed from different sources.
@@ 8 Conclusions
In this paper, by uncovering the impact of the newly emerged physical mechanism, formed as a result of halting BHL accretion around a Kerr black hole, on the generation of QPOs, we link plasma dynamics to both LFQPOs and HFQPOs within a unified framework. To achieve this goal, we model the complete disappearance of the shock cone formed around the black hole. The shock cone is a physical mechanism previously established in strong gravitational fields around the black holes using the BHL mechanism in both general relativity and alternative gravity theories. We examine how this shock cone mechanism evolves after the cessation of BHL accretion and observe the formation process of the newly emerging plasma structure. Eventually, after a long dynamical evolution, the matter in plasma reaches a steady-state configuration around the black hole, forming a quasi-periodically oscillating plasma structure. We investigate how the dynamic evolution of the plasma structure and the resulting precession motion depend on the black hole spin parameter and asymptotic velocity. Finally, by performing a power spectrum analysis, we identify the QPO frequencies that emerge from the system.
@@
It has been observed that the asymptotic velocity used in modeling the shock cone around a black hole via BHL accretion affects the plasma structure formed after BHL is halted, and consequently, the resulting QPOs. For a rapidly rotating black hole with $a/M=0.9$ and an asymptotic velocity of $V_{\infty}/c=0.1$, meaning that the flow is sonic, LFQPOs are not observed. However, LFQPOs appeared at other asymptotic velocities. Therefore, for LFQPOs to be observable, the Lense-Thirring effect, which arises solely from the black hole’s rapid rotation and its warping of spacetime, would not be sufficient. By contrast, the literature notes that different mechanisms around black holes, for instance, disk truncation (Ingram2009MNRAS, ), have demonstrated that LFQPOs can also arise in subsonic flows. Thus, our results should be regarded as complementary rather than exclusive. Additionally, it is known that indirectly the velocity of the gas and, consequently, its angular momentum contribute to the formation and regulation of QPOs Mukhopadhyay_2009 ; 2009MNRAS.398.1105R . However, based on the numerical simulations, it has been found for the first time that if the plasma exhibits supersonic behavior, LFQPOs emerge.
@@
As seen in numerical simulations and Table 2, in addition to the fundamental HFQPOs, other HFQPOs have been observed that generate different resonance modes (e.g. $3:2$ and $2:1$), which are consistent with observations. These frequencies result from the fundamental mode itself and the nonlinear couplings that form between subsequently generated QPOs. Furthermore, as shown in Table 2, these couplings occur only when the black hole is rapidly rotating. Thus, the presence of nonlinear couplings in rapidly rotating black holes suggests that additional oscillatory modes are introduced due to the strong frame-dragging effect and the differential rotation of the accretion disk.
The results obtained here show consistency with the observations. In microquasars $GRS1915+105$, XTEJ1550564XTEJ1550-564, and GROJ165540GROJ1655-40, the observed QPO frequencies range between 3020030-200 Hz. The numerical simulations presented in this study yield HFQPO frequencies in the range of 364936-49 Hz, which are consistent with black holes of approximately stellar mass $M\sim 10M_{\odot}$. Additionally, the fact that the numerical simulations show double and triple the fundamental frequency suggests that the $3:2$ and $2:1$ resonance states observed in black hole QPOs can be explained through these results. This supports the idea that QPOs in black hole systems may arise from parametric resonances in the accretion disk. On the other hand, the observed frequency range for LFQPOs in black hole $X$-ray binaries is 0.1300.1–30 Hz, whereas the numerical simulations in this study produced LFQPO frequencies in the range of 1.522.61.5-22.6 Hz, which aligns well with observations. However, LFQPOs did not appear when $V_{\infty}/c=0.1$ for $a/M=0.9$. This indicates that LFQPOs depend on the accretion state, reinforcing their consistency with observational data.
@@
We will continue working in the future to uncover the effects of the black hole spin parameter, accretion dynamics, and the physical parameters governing these processes, as well as the behavior of plasma in a strong gravitational field. In this study, magnetic fields and vertical instabilities are neglected; however, in future work, we plan to incorporate modified gravity theories, magnetic fields, and full 3D simulations to examine how these additional physical effects influence the QPO frequencies identified here. We will utilize modified gravity theories to investigate the influence of these alternative gravitational models on the QPO frequencies that we have identified in this study. Through this approach, we aim to determine the impact of the scalar field, which represents the existence of dark matter, on the results obtained here.